Two trains leave the same station heading in opposite directions, each at 10 miles per hour. After 1/2 hour they each turn left and continue on at 5 miles per hour for 2 hours. At that point how far apart are they?
Answer by checkley75(3666)

WHAT YOU HAVE ARE 2 RIGHT TRIANGLES WITH ONE SIDE 5 MILES THE OTHER SIDE 10 MILES. NOW WE NEED TO FIND THE THIRD SIDE (HYPOTENUSE) THUS:
5^2+10^2=H^2
25+100=H^2
125=H^2
H=SQRT125
H=11.28 MILES. NOW WE HAVE TO DOUBLE THIS VALUE SO 2*11.28=22.36 MILES APART ARE THESE 2 TRAINS.